OpenStudy (anonymous):

please help the set of integers as Z−....−3,−2,−1,0,1,2,3.... the greatest integer of x is denoted as ⌊x⌋ is defined to be the largest interger k with k≤x discuss with your classmates how ⌊x⌋ may be described as a piecewise defined function

OpenStudy (anonymous):

also is ⌊a+b⌋=⌊a⌋+⌊b⌋ always true? what if a or b was an integer? test some values make a conjecture and explain your result

OpenStudy (anonymous):

This is the floor function. It would truncate the number down to an integer value ot basically remove the digits after the decimal point. This function would give a constant value between integer values. For example: f(x)=0 for 0=<x<1 f(x)=1 for 1=<x<2 and so on. f(x)+f(y) would not always be equal to f(x+y), consider 1.4 and 1.7 as x and y. However f(x)+f(y)=f(x+y) if x-f(x)+y-f(y)<1. And that applies to integer values. Let me know if you need more explanation.